# Multiplexed Processing of Quantum Information Across an Ultra-wide Optical Bandwidth

### v3

@@ Abstract
Multi-channel quantum teleportation, which we propose and analyze here, is another example for harnessing the optical bandwidth to multiplex an important protocol of quantum information. Again we note that this protocol should not be judged as immediately applicable for technology, but rather as an illustration of the range of possibilities that our multiplexing scheme offers for quantum information. This new protocol is a broadband, multiplexed version of the Braunstein’s & Kimble’s protocol (suggested in [26] and demonstrated in [27]). For this scheme we utilize two sources of broadband squeezed vacuum (OPAs) to simultaneously teleport a set of quantum states across the spectrum of a general broadband field with arbitrary two-mode quadratures at each frequency (see figure 3).
As opposed to the QKD protocol above, which operates in the regime of low squeezing with single pairs of entangled photons, the teleportation protocol we now discuss operates best at the high-squeezing regime. Thus, for the sake of presentation simplicity only, let us assume initially that the two sources are ”infinitely squeezed” to the level that we can completely neglect one of their quadratures. Later we will alleviate this assumption and consider the implications of finite squeezing to the teleportation precision. When the two highly squeezed sources (marked (1) and (2) on the figure) are mixed on a beam splitter (BS) with the correct phase, they generate an entangled quantum state, where the quadratures at the two outputs of the BS (marked (3) and (4)) are quantum-correlated. To teleport the input state, we mix it with one of the entangled arms (on another BS) and measure quadratures of the BS outputs (marked (5) and (6)). Based on this measurement, we introduce a quadrature shift (marked (7)) to the unmeasured entangled arm (marked (4)) to reproduce the original quadratures of the input state at the teleportation output (marked (8)).
To describe the steps of this protocol, we will use the field operator $\hat{a}_{\omega}$ at each frequency which can be decomposed into quadratures as $\hat{a}_{\omega}=\hat{x}_{\omega}+i\hat{y}_{\omega}^{\dagger}$ (here we use the definition of the two-mode quadratures, $\hat{x}_{\omega}\equiv\frac{1}{2}\left(\hat{a}_{\omega}+\hat{a}_{-\omega}^{\dagger}\right)$ and $\hat{y}_{\omega}\equiv\frac{1}{2i}\left(\hat{a}_{\omega}-\hat{a}_{-\omega}^{\dagger}\right)$ [28, 29], which is a convenient generalization of the standard single-mode quadratures to the two-mode squeezed pair of signal-idler modes). With this definition, we can represent the input field operator as $\hat{a}_{\omega,in}=\xi(\omega)\hat{x}_{\omega}+i\eta(\omega)\hat{y}_{\omega}$ and the field operator of the squeezed state generated by an OPA as $\hat{a}_{\omega,OPA}=X(\omega)\hat{x}_{\omega}+iy(\omega)\hat{y}_{\omega}$, where without loss of generality, $X$ represent the stretched two-mode quadrature (at each frequency $\omega$), whereas $y$ represent the squeezing of the other quadrature (dictated by the parametric gain $g(\omega)$ of the squeezers at each frequency, which ideally set $X\!=\!e^{g},y\!=\!e^{-g}$). For convenience, we will drop the frequency index from now on (assuming we look at a specific frequency component of the broad spectrum). Let us now describe in detail each step of this protocol:
@@
    The quadratures of the two encoded states are measured with parametric homodyne measurement simultaneously across the two-mode spectrum, such that the quadrature $\hat{x}$ is measured for $\hat{a}_{5}$ and $\hat{y}^{\dagger}$ for $\hat{a}_{6}$. As a result, we obtain information on the difference of the signals quadrature, $\frac{1}{\sqrt{2}}\left(\xi-X\right)$ and $\frac{1}{\sqrt{2}}\left(\eta-Y\right)$ without any knowledge about the quadratures themselves.
@@
Figure 3: The multiplexed quantum teleportation protocol. Note that in the figure dashed lines represent classical signals, electrical or optical whereas the solid lines are quantum optical signals. Two OPAs are generating two broadband orthogonal squeezed states, squeezed in the $\hat{x}$ and $\hat{y}$ directions (1 and 2). These states are interfered using a beam splitter to produce two quadrature-entangled states (3 and 4). Half of the entangled state (4) is interfered with an input signal (in) using a beam splitter and produces two new interference states (5 and 6). Theses states are sent to two orthogonal homodyne detectors, which measure the quadratures $\hat{x}$ and $\hat{y}^{\dagger}$ respectively to obtain the difference between the entangled state quadratures (3) and the input signal (in). Finally, a shift controlled by the measured quadratures is performed on the rest of the entangled state (3) (using a spectrally shaped classical source) to reproduce the original input state at the output (8).
@@
indicating that the residue of the squeezed quadratures act as a source of noise, added to the teleportation output. Thus, in order to reduce these errors, it is important to maximize the squeezing of the original signals. For example, one can enhance the squeezing level by replacing the single-pass OPAs in our protocol with multi-pass OPOs (optical parametric oscillator) that offer higher squeezing (up to 15 dB demonstrated [30]).
.3 Experimental Demonstration
Figure 4: Proof-of-principle demonstration of multiplexed QKD: (a) The experimental configuration. The pump beam (purple) was generated by a CW, single-frequency laser at $780nm$ of $\sim 1-4W$ output power. The two OPAs of the SU1,1 interferometer were realized in a single periodically polled LiNbO3 crystal (PPLN in the figure), by passing through the PPLN medium twice: forward for generation of the SPDC photons across a wide spectrum of signal and idler (red and green); and backwards, for parametric homodyne detection. Alice encodes the SPDC light using a spectral shaper for the SPDC and a phase modulator (EOM) for the pump, which is separated off from the SPDC by a harmonic separator (H.S.). To encode information, Alice employs a spectral shaper that consists of a grating, a lens and a spatial light modulator (1x12,288 liquid crystal array SLM). The EOM stabilizes the phase of the pump to that of the SPDC, compensating for phase noise and drifts, by using an active feedback loop (not shown). The EOM also allows Bob to select the measurement basis (same for all channels across the spectrum). Finally, the three beams are reflected back into the OPA crystal, which now acts as Bob’s detector. The reflected SPDC light is separated from the pump with another harmonic separator mirror (H.S.) and the resulting SPDC spectrum is recorded on a home-built spectrometer that consists of a grating, a lens and a Linear CCD camera. (b) Experimental results: the normalized spectral intensities of all channels are shown for measurement in basis ”1” (top) and ”2” (bottom). The colors represents the encoded phase by Alice of each channel, and the height is the normalized intensity (detection probability) measured by Bob. Evidently, Bob can easily decode only the channels that were measured at the correct basis.
@@
During the experiment the pump passes through the OPA and generates a broadband spectrum of signal and idler pairs across ${}150nm$ bandwidth around $1560nm$. The pairs are separated from the pump with a Harmonic Separator (H.S) and sent to a spectral shaper to encode Alice’s information by modulating the phase of each frequency pair (channel) to $0,\pi$ or $\frac{\pi}{2},\frac{3\pi}{2}$ randomly (step 1). Simultaneously, the pump passes through an EOM used to stabilize the phase of the pump to that of the pairs, as well as to choose the measurement basis for Bob (same basis for all channels in this case). Finally, the pump and the pairs are reflected back to recombine and pass through the OPA once more in the opposite direction, completing the SU1,1 interference. To implement Bob’s parallel detection of all channels we measure the output spectral intensity using a home-built spectrometer composed of a grating and a line CCD-camera (step 2).
@@
To maximize the data capacity while preserving the security of the protocol, the spectral width of the channels at the spectral shaper was chosen to be the smallest possible without leaking to the neighboring channels (see experimental verification in the methods, section I.2). This allowed us to encode and decode in this preliminary configuration up to 23 channels in parallel with a bare interference contrast of $75\%$. This 23-fold enhancement of the data capacity (compared to a single channel) is far from any fundamental limit. It can easily be pushed up to $>\!200$ channels by improving the spectral resolution of both the encoding shaper and the spectrometer with standard technology of optical wavelength division multiplexing (see section II.1 in the supplementary).
@@
As discussed above, our scheme is a CV, multiplexed version of the well-known BB84 protocol, indicating that all the security proofs of BB84 are directly applicable to our protocol as well. We chose to demonstrate the immunity of our scheme to the steal-attack, i.e. an attempt of Eve to split off some of the light between Alice and Bob, which is a most common attack on QKD. If Eve uses a beam splitter to ”steal” part of the quantum state, she inevitably will reduce the interference contrast for Bob. We therefore simulated Eve’s operation by introducing loss at the spectral shaper. Although the initial loss in our simple configuration was relatively high (${}44\%$, probably due to imperfect components and alignment), we could still clearly detect even a small additional loss of $5\%$ since it visibly reduces the contrast (compared to the measurement error). For more details about the loss detection see the methods, section I.3.
.4 Discussion
High data throughput is an important attribute of any information processing scheme. Although the bandwidth of standard broadband sources of squeezed light and entangled photons can easily exceed $10$THz (even up to an octave in frequency [21]), the bandwidth resource of the light is yet to be utilized, mainly due to the lack of efficient measurement techniques with sufficient bandwidth. Our method harnessed optical parametric homodyne in order to efficiently utilize the optical bandwidth to increase the processing capacity by several orders of magnitude. With this method we proposed and demonstrated a multiplexed, BB84-like protocol of QKD, as well as a new quantum teleportation protocol. In spite of the conceptual difference between the teleportation and the QKD applications, they both shared the same set of tools for broadband state generation, broadband state manipulation and broadband state measurement.
If we examine these quantum tools in a broader view, we can realize that they are applicable in the general context of broadband quantum information processing, well beyond QKD and teleportation. An evident example will be to realize frequency multiplexed versions of other protocols of quantum communication, such entanglement-based QKD [3], quantum coin flipping [1], entanglement based sensing [7], etc.. Furthermore, repeating those operations and combining them in various manners is key to implement a broadband quantum network with ultra-fast communication speed.
Another thrust of possible application is towards high-bandwidth quantum computation. Specifically, the ability to simultaneously generate, control and measure a large set of separated squeezed qubits ( implemented as signal-idler pairs, as presented above), along with a multiplexed two-qubit operation, is sufficient for universal quantum computation that will exploit the frequency dimension to generate much larger entangled states with the same squeezing resources [31, 32, 33]. Note that a multiplexed two-qubit gate was already demonstrated across the quantum frequency comb of a broadband OPO, that is either pumped by several frequencies [31] or phase-modulated in time [34]. Such a multiplexed quantum computer will be naturally compatible with the multiplexed quantum network, mentioned above.
.5 Acknowledgements
@@
This research was funded in part by SPARQL consortium, under the QuantERA program of the EU.

### v4

@@ 3 Multiplexed Quantum Teleportation
Multi-channel quantum teleportation, which we propose and analyze here, is another example for harnessing the optical bandwidth to multiplex an important protocol of quantum information. Again we note that this protocol should not be judged as immediately applicable for technology, but rather as an illustration of the range of possibilities that our multiplexing scheme offers for quantum information. This new protocol is a broadband, multiplexed version of the Braunstein’s & Kimble’s protocol (suggested in [32] and demonstrated in [33]). For this scheme we utilize two sources of broadband two-mode squeezed vacuum (OPAs) to simultaneously teleport a set of arbitrary quantum states across the spectrum of a general broadband field (see figure 3).
Before delving into the details of the teleportation protocol, let us briefly review the continuous-variable description of quantum optics with broadband fields using two-mode quadratures (rather than the discrete basis of photon number). We adopt the formalism of two-mode quadratures $\hat{x}_{\omega}\equiv\frac{1}{2}\left(\hat{a}_{\omega}+\hat{a}_{-\omega}^{\dagger}\right)$ and $\hat{y}_{\omega}\equiv\frac{i}{2}\left(\hat{a}_{\omega}^{\dagger}-\hat{a}_{-\omega}\right)$ [29, 34], where the $\omega,-\omega$ indices denote the signal and idler frequencies relative the the center carrier frequency of $\omega_{p}/2$ (later to become independent spectral channels of teleportation). The two mode quadratures are a generalization of the well known single-mode quadratures that enables to treat squeezing on equal footings, be it degenerate and non-degenerate (two-mode) squeezing under a unified framework (maintaining the canonical commutation $\left[\hat{x}_{\omega},\hat{y}_{\omega}\right]\!=\!\frac{i}{2}$) .
With two-mode quadratures we can decompose the field operator $\hat{a}_{\omega}$ at any frequency $\omega$ as $\hat{a}_{\omega}\!=\!\hat{x}_{\omega}+i\hat{y}_{\omega}^{\dagger}$ Adopting the Heisenberg picture of quantum dynamics, the field operators evolve in time under the Hamiltonian of the system. Thus, the field operator of a squeezed state generated by an OPA is $\hat{a}_{\omega,OPA}=e^{g}\hat{x}_{\omega}+ie^{-g}\hat{y}_{\omega}^{\dagger}=X(\omega)\hat{x}_{\omega}+iy(\omega)\hat{y}_{\omega}^{\dagger}$, where $X\!=\!e^{g}$ represents the stretched two-mode quadrature and $y\!=\!e^{-g}$ the squeezed ($g(\omega)$ is the parametric gain). Similarly, we can represent the field operator of the input state of channel $\omega$ (to be teleport) as $\hat{a}_{\omega,in}=\xi(\omega)\hat{x}_{\omega}+i\eta(\omega)\hat{y}_{\omega}^{\dagger}$, where the teleportation task is to transfer both quadratures $\xi(\omega),\eta(\omega)$ from the input to the output without measuring any one of them. For convenience, we will drop the frequency index from now on (assuming we look at a specific frequency if component of the broad spectrum).
As opposed to the QKD protocol above, which operates in the regime of low parametric gain, where the squeezing is negligible and the generated light can be considered as a random stream of single bi-photons, the teleportation protocol here operates best at the high parametric gain regime of multiple photons and high-squeezing. Let us therefore assume (for simplicity of the presentation) that the two sources are "infinitely squeezed", such that we can completely ignore the squeezed quadratures. Later we will alleviate this assumption and consider the implications of finite squeezing to the teleportation precision.
Let us now describe in detail each step of this protocol, which is also visually presented in figure 3:
@@
    The quadratures of the two encoded states are measured (intensity and phase) with parametric homodyne measurement[21, 35, 36, 37] simultaneously across the two-mode spectrum, such that the quadrature $\hat{x}$ is measured for $\hat{a}_{5}$ and $\hat{y}^{\dagger}$ for $\hat{a}_{6}$. As a result, we obtain information on the difference of the signals quadrature, $\frac{1}{\sqrt{2}}\left(\xi-X\right)$ and $\frac{1}{\sqrt{2}}\left(\eta-Y\right)$ without any knowledge about the quadratures themselves.
@@
Figure 3: The multiplexed quantum teleportation protocol. Two OPAs are generating two broadband orthogonal squeezed states, one in the $\hat{x}$, the other in $\hat{y}$ directions ($X,iY$ marked (1) and(2)). These states are interfered using a beam splitter to produce two quadrature-entangled states ($X\pm iY$ marked (3 and (4)). One of the entangled states (3) is sent on to the teleporting site, while the other (4) is mixed on a beam splitter (BS) with the input signal to be teleported ($\xi+i\eta$, marked (in)), which produces two interference outputs (marked (5) and (6)) that are them measuredby two orthogonal broadband parametric homodyne detectors - one to measure the $\hat{x}_{\omega}$ quadrature and the other $\hat{y}^{\dagger}_{\omega}$ across the spectrum, which provides the instantaneous difference between the entangled state quadratures $\xi-X,\eta-Y$ without measuring $\xi,X,\eta,Y$ themselves. Finally, these measurement results are communicated classically (over a classical channel, marked thick black) to the teleporting site, where it is used to generate a coherent classical local oscillator $\alpha_{\text{shift}}(\omega)\sim\xi-X+i\eta-Y$ (using a spectrally shaped broadband classical source / ultrashort pulse). This LO $\alpha_{\text{shift}}(\omega)$ is then mixed on an unbalanced beam-splitter with the received state (3) in order to shift it by the exact amount needed to reproduce the input state $\xi+i\eta$ at the output (8) on every frequency channel.
@@
indicating that the residue of the squeezed quadratures acts as a source of noise, added to the teleportation output. Thus, in order to reduce these errors, it is important to maximize the squeezing of the original signals and to minimize the loss of the transmitted quantum state, $\hat{a}_{3}$ (as loss reduces squeezing and introduces vacuum noise). For example, one can enhance the squeezing level by replacing the single-pass OPAs in our protocol with multi-pass OPOs (optical parametric oscillator) that offer higher squeezing (up to 15 dB demonstrated [38]).
Just like the QKD protocol above, the teleportation requires a shared classical resource - a phase reference. While one possible implementation is to send the pump along with the squeezed spectrum, many other options can be conceived according to the technical limitations of the classical communication. We highlight other possibilities in the discussion below.
4 Experimental Demonstration
@@
Figure 4: Proof-of-principle demonstration of multiplexed QKD: (a) Experimental configuration The pump (purple) was a CW, single-frequency laser at $780nm$ of $\sim 1-4W$ power. The two OPAs of the SU1,1 interferometer were realized in a single periodically polled LiNbO3 crystal (PPLN), by passing through the crystal twice: forward for generation of the broadband bi-photons (red and green); and backwards, for parametric homodyne detection. Alice encodes the bi-photons spectrum using a spectral shaper and a phase modulator (EOM) for the pump, which is separated off from the bi-photons by a harmonic separator (H.S). Alice encodes information with a spectral shaper that consists of a grating, a lens and a spatial light modulator (SLM - 1x12,288 linear array of liquid crystal pixels). The EOM stabilizes the phase of the pump to that of the bi-photons, compensating for phase noise and drifts with an active feedback loop (not shown). The EOM also allows Bob to select the measurement basis (same for all channels across the spectrum). Finally, the three beams are reflected back into the OPA crystal, which now acts as Bob’s detector. The output bi-photons beam is separated from the pump again with another H.S mirror and the resulting bi-photons spectrum is measured on a home-built spectrometer that consists of a grating, a lens and a Linear CCD camera. (b) Experimental results: the normalized spectral intensities of all channels, which represent the detection probability of a photon per channel, are shown for measurement in basis "1" (top) and "2" (bottom). The colors represents the encoded phase by Alice for each channel, and the height is the normalized intensity (detection probability) measured by Bob. Evidently, Bob can easily decode only the channels that were measured at the correct basis.
@@
During the experiment the pump passes through the OPA and generates a broadband spectrum of signal and idler pairs across ${}150nm$ bandwidth around $1560nm$. The pairs are separated from the pump with a Harmonic Separator (H.S) and sent to a spectral shaper to encode Alice’s information by modulating the phase of each frequency pair (channel) to $0,\pi$ or $\frac{\pi}{2},\frac{3\pi}{2}$ randomly (step 1). The spectral shaper also serves to compensate for the chromatic dispersion that the signal and idler accumulate during the propagation through the optical configuration between the OPAs. Simultaneously, the pump passes through a phase modulator (EOM) used to stabilize the phase of the pump relative to that of the pairs, as well as to choose the measurement basis for Bob (same basis for all channels in this case). Finally, the pump and the pairs are reflected back to recombine and pass through the OPA once more in the opposite direction, completing the SU(1,1) interference. To implement Bob’s parallel detection of all channels we measure the output spectral intensity using a home-built spectrometer composed of a grating and a line CCD-camera (step 2). We tuned Bob’s parametric gain to roughly equal that of Alice by polarization manipulations on the pump beam (not shown). Practically, this could also be achieved by classical amplification, since the gain (pump) is a classical resource.
@@
To maximize the data capacity while preserving the security of the protocol, the spectral width of the channels at the spectral shaper was chosen to be the smallest possible without leaking to the neighboring channels (see experimental verification in supplement 1). This allowed us to encode and decode in this preliminary configuration up to 23 channels in parallel with a bare interference contrast of $75\%$. This 23-fold enhancement of the data capacity (compared to a single channel) in this proof-of-principle experiment is far from any fundamental limit for two reasons. First, the number of channels can easily be pushed up to $>\!200$ channels by improving the spectral resolution of both the encoding shaper and the spectrometer with standard technology of optical wavelength division multiplexing (see supplement 1). Second, the low communication speed of the single channel ( $\sim\!100$Hz due to the slow speed of the liquid crystal SLM) can be drastically uplifted with real-time phase modulators (electro-optic or acousto-optic), which can reach up to GHz rates.
For QKD to operate securely, it is preferable to maintain a low parametric gain with $N_{q}\ll 1$ photons per quantum state. The bandwidth of each channel in our experiment was $\sim 100GHz$, which indicates a coherence time of $\tau_{c}\sim\frac{1}{100GHz}=10ps$ per channel, which was much longer than the phase delay induced by the SLM (of order 1 optical cycle, maximum), certifying ideal visibility [39, 40]. The pump power in our experiment ($4W$) generated a flux of up to $10^{7}$ photons/sec per channel (as measured on our CCD), which indicates $N_{q}\leq 10^{-4}$ per quantum state - safely within the low gain regime. The integration time per bit in the QKD protocol can then be chosen at $10^{4}\tau_{c}=100ns$, such that the average detection probability of a photon will be near $N_{detect}\sim 1$ per bit. To demonstrate that the interference outcome is easily identified in our experiment for the different bases, we purposely used a much longer integration time on our CCD (about $1ms$) with $\sim 10^{4}$ photo-detection events per integration time, where the measured intensity directly represents the probability of detection for every channel.
@@
As discussed above, our scheme is a CV, multiplexed version of the well-known BB84 protocol, indicating that all the security proofs of BB84 are directly applicable to our protocol as well. We chose to demonstrate the immunity of our scheme to the steal-attack, i.e. an attempt of Eve to split off some of the light between Alice and Bob, which is a most common attack on QKD. If Eve uses a beam splitter to "steal" part of the quantum state, she inevitably will reduce the interference contrast for Bob. We therefore simulated Eve’s operation by introducing loss at the spectral shaper. Although the initial loss in our simple configuration was relatively high (${}44\%$, probably due to imperfect components and alignment), we could still clearly detect even a small additional loss of $5\%$ since it visibly reduces the contrast (compared to the measurement error). For more details about the loss detection see supplement 1.
5 Discussion
To perform quantum information processing with our tools, some classical and technical resources are required, such as the classical phase reference and the spectral phase modulation device. In the description of our protocols above and in our experimental demonstration we chose simple implementations of these resources: sending the pump as a phase reference and using a Fourier domain shaper as the broadband phase modulator. However, many other implementations are possible that may be advantageous for specific realizations under given technical constraints.
For example, the update rate of the liquid-crystal based SLM in our spectral shaper is rather slow ($\sim 100$Hz), which would limit the data rate per channel to be impractically low for a realistic QKD application. In addition, the transmission of the fiber connecting the communicating parties may be too lossy in a real scenario to transmit the pump laser from Alice to Bob; and the chromatic dispersion of that fiber may require handling / compensation that is beyond the capability and dynamic range of the spectral shaper.
While these technical considerations are very important for a practical realization, they are classical constraints that can be overcome in various classical means, depending on the technical details of the implementation. For example, instead of the pump one can utilize the degenerate frequency channel at $\omega_{p}/2$ as phase reference (the center of the SPDC spectrum) to send a laser as a classical local oscillator that is less attenuated in the connecting fiber (and can be classically amplified, if needed). Other methods for distributing a time-base across the fiber can be adequate as well [41, 42, 43]. Fast spectral modulation can be implemented using arrays of electro-optic modulators [44] or fast acoustic modulators [45], or even internally in the fiber with a space-division-multiplexing networks [46]. Chromatic dispersion can be compensated either by spectral modulators (as we did in our experiment above) or by incorporating a dispersion compensating fiber.
6 Conclusion
We presented a set of tools to efficiently multiplex quantum information processing across a broad bandwidth of squeezed light with a large number of frequency channels. To demonstrate the capabilities of these tools, we developed multiplexed versions for two of the most important quantum communication protocols: QKD, which was also demonstrated in a proof-of-principle experiment on 23 parallel channels and quantum teleportation. Since broadband sources of squeezed light and entangled photons can easily exceed $10$THz (up to an octave in frequency [18]), our new method can potentially speed-up quantum processing by orders of magnitude compared to the standard, single channel implementations.
In a broader view, our set of tools is applicable to quantum processing in general, and can be used as building blocks for frequency multiplexed versions of many other quantum protocols, such as quantum coin flipping [1], entanglement-based QKD [3] and entanglement based sensing [7], which we plan to pursue in the future. Furthermore, our tools can be extended to include two-qubit operations (that have already been demonstrated across the quantum frequency comb of a broadband OPO [47, 48]), which can pave the way towards high-bandwidth quantum computation with the potential to be fast, scalable (in the number of qubits) and compatible with broadband quantum communication networks [47, 49, 50].
[…]
@@ 6 Conclusion
Acknowledgements This research was funded in part by SPARQL consortium, under the QuantERA program of the EU.
[…]
Supplemental document See Supplementary Material for supporting content.